The distance light travels in one year is approximately 5,870,000,000,000 miles. The distance light travels in 100 years is: (a) 587 × 108 miles, (b) 587 × 1010 miles, (c) 587 × 10-10 miles, (d) 587 × 1012 miles, (e) 587 × 10-12 miles

Grade:12th pass

2 Answers

User
7 years ago
The distance of the light travels in 100 years is: 

5,870,000,000,000 × 100 miles.

= 587,000,000,000,000 miles.

= 587 × 1012 miles.

Answer: (d)
MOHIT GARG
7 years ago
DISTANCE COVERED IN I YEAR-587*10^10 MILES
THUS DISTANCE COVERED IN 100 YEARS=597*10^12*100=
587*10^12 MILES
BRO ALWAYS RIGHT THINGS IN SCIENTIFIC WAY TO SOLVE EASILY